12(4c-5)=7c-3
x/4+1=x/5+2
-5x/-3-1/4=-x/-6-1/8

Answer by Boreal(15235)

12(4c-5)=7c-3 ; 48c-60=7c-3 ; 41c = 57 c= 57/41

x/4+1=x/5+2
Multiply everything by 20; 5x+20=4x+40; x=-20
-5x/-3-1/4=-x/-6-1/8
(5x/3)+(1/8)+(x/6)=(1/4) ; I like dealing with positive signs
Multiply everything by 24
40x + 3 +4x=6
44x=3
x=(3/44)